working together,Carmel and Menchie can encode a document in 3 hours.Working alone,Menchie can do the job in 7 hours.How long does it take carmel to do the job working alone?
Answer by Grinnell(63)

let m be the time Menchie takes to do the job alone, which is 7 hours
let c be the time Carmel can do the job alone. Let's go!!
1/c+1/7=1/3, that is, working together they can do the job in three hours,
this is the part of the job that can be done PER HOUR.
Let's do the algebra (multiply thru by 21c!)
21+3c=7c
21=4c
5.25=c
(1/4 of an hour is 15 minutes, no problem)
5 hours and 15 minutes is the time Carmel can do the job alone.